What is the escape velocity at the surface of the sun?


1
Expert's answer
2021-10-20T10:43:07-0400

The escape velocity is


"v=\sqrt\frac{2GM}{R}=\sqrt\frac{2(6.673·10^{-11})(1.989 · 10^{30})}{6.957·10^8}=618\text{ km/s}."